10987654321:

After completing the fraction division 8 divided by StartFraction 1 Over 6 EndFraction, Patel used the multiplication below to check his work. StartFraction 1 Over 48 EndFraction times 6 = StartFraction 1 Over 48 EndFraction times StartFraction 6 Over 1 EndFraction = StartFraction 6 Over 48 EndFraction = StartFraction 1 Over 8 EndFraction Which is the most accurate description of Patel’s work? Patel found the correct quotient and checked his work using multiplication correctly. Patel found the correct quotient but checked his work using multiplication incorrectly. Patel found an incorrect quotient but checked his work using multiplication correctly. Patel found an incorrect quotient and checked his work using multiplication incorrectly.

lilianamendez2:

StartFraction 1 Over 48 EndFraction times 6 = StartFraction 1 Over 48 EndFraction times StartFraction 6 Over 1 EndFraction = StartFraction 6 Over 48 EndFraction = StartFraction 1 Over 8 EndFraction = \(\LARGE \frac{1}{48} \times 6 = \frac{1}{48} \times \frac61=\frac{6}{48}=\frac{1}{18}\) ???